# Secure Outsourced Decryption for HE-based Privacy-preserving Cloud Computing System

### v1

@@ V Security Analysis
We initially present formal proof of the security of OstDec in Theorem V.1, reducing the protocol’s security to the RLWE and NTRU assumptions. This stems from the fact that our blinded secret key can be viewed as a $(\gamma_{s},\gamma_{t},q)$-NTRU instance given in Definition V.1. Recovering the secret key from the blinded one corresponds to solving the $(\gamma_{2},\gamma_{t},q)$-search NTRU problem in Definition V.2.
@@
The cryptosystem consisting of an underlying RLWE-based HE scheme and the outsourcing decryption protocol of direct connection mode is semantically secure, provided that the RLWE and NTRU-search assumptions hold where RLWE instances and an NTRU instance are public together.
@@
    • $pk,s\leftarrow HE.KeyGen(1^{d})$,
    • $(t,t^{-1})\leftarrow OtsDec.SkBdKeyGen()$,
    • $\tilde{s}\leftarrow OtsDec.SkBd(s,t^{-1})$.
@@
    $\mathcal{U}_{H}^{\prime}$ computes $c=HE.Enc(pk,m)$ where $m$ is sampled uniformly from the plaintext space.
@@
[49] A $(\gamma_{s},\gamma_{t},q)$-NTRU instance is $r\in R_{q}=Z_{q}[X]/(X^{n}+1)$ such that (i) $r=s\cdot t^{-1}\mod{q}$ and (ii) $||s||_{2}\leq\frac{\sqrt{q}}{\gamma_{1}},||t||_{2}\leq\frac{\sqrt{q}}{\gamma_{2}}$
@@
[49] The $(\gamma_{2},\gamma_{t},q)$-search NTRU problem asks, given a $(\gamma_{s},\gamma_{t},q)$-NTRU instance $r$, to recover $(s,t)\in R^{2}$ such that (i) $r=s\cdot t^{-1}\mod{q}$ and (ii) $||s||_{2}\leq\frac{\sqrt{q}}{\gamma_{1}},||t||_{2}\leq\frac{\sqrt{q}}{\gamma_{2}}$
@@ V-A Attacks on the secret key blinding technique
Given a blinded private key $\tilde{s}=st^{-1}$, along with public parameters $\{d,q,h\}$ for the blinding algorithm, the adversary’s goal is to recover either $s$ or $t$. $h$ is not necessarily a public parameter, but we assume it for the simplicity of security discussion. We denote the distributions of $s$ and $t$ as $\chi_{sk}$ and $\chi_{t}$, respectively. We also denote the set of all elements that $s$ can take as $S_{sk}$ and the security parameter as $\lambda$.
@@
TABLE I: Lattice Reduction for $L(\tilde{s},\alpha)$ with configured by $h=5$ and $\log{q}=33$
@@
The probability for us to guess a correct set of $J$ for $t$ or any of its shifted versions is computed with the following equation proposed by xxx:
@@
Taking the result of Equation 10 as $p$, the expected time for a zero-forced attack with specific $r$ to find $\tau$ or any of its shifting versions in some $L(\alpha,\tilde{s})^{ZF}$ is computed by $1/p\cdot\text{oprs}^{ZF}$, where $\text{oprs}^{ZF}$ denotes the number of operations required for reducing one $L(\alpha,\tilde{s})^{ZF}$ instance. Table 3 demonstrates the probability $p$ and the zero-forced attack complexity as $r$ increases with different choices of $h$. The complexity of the attack corresponds to the number of elementary operations in one BKZ iteration[12, 58]: $\log{(\text{oprs})}=poly(\beta)+\log{dim}+7$, where $poly(\beta)=0.000784314\beta^{2}+0.366178\beta-6.125$.
@@ V-B Security Parameterization
Based on the attacks above against OstDec security, we can parameterize the security of the OstDec protocol. We consider three security parameter cases: $\lambda=128,192,256$. For each $\lambda$, we discuss the required values of $h$ and $q$ when $d=13,14,15$, and $16$. During the parameterization, $h$ is selected such that a zero-forced attack requires a complexity no smaller than $2^{\lambda}$ elementary operations. For a given $h$, we fix the sampling method of $t$ as $t\leftarrow\Pi_{\text{SkBdKeyGen}}(R_{q},h_{1})\cdot(\Pi_{\text{SkBdKeyGen}}(R_{q_{2}},h_{2}))\mod{q}$, where $h_{1}\cdot h_{2}-\text{min}(h_{1},h_{2})\geq h$. The sampling space determined by $q$ and $h$ should be large enough to guard against brute-force attacks. For adversaries enumerating the $t$ using the sampling method above, $q$ and $h$ should satisfy the following equation:
@@
TABLE III: parameterization for OtsDec protocol
@@ VI Implementation
After all the discussion regarding the security of our proposed outsourced decryption protocol, we now move on to the implementation.
Figure 4: Local decryption time with different polynomial degree and fixed $\log{q}\leq 64$)
Figure 5: Local decryption time with different length of coefficient modulus and fixed $d=2^{15}$
@@ VI-A Efficiency
To evaluate the protocol’s performance on practical RLWE-based HE parameter sets, we compare the local decryption time of the original decryption and our outsourced method under parameter sets with $d={13,14,15,16}$ (see Table xxx for their security levels corresponding to the RLWE assumption). The comparison results are presented in Figures 4 and 5.
@@
It can be observed that the differences between the curves for various $h$ are minimal, as in the sampling method of $t$, with $h_{1}$ fixed at $6$, $h_{2}$ only needs to be within the range $[3,8]$ to meet all three security categories. On the other hand, the time taken by the original decryption algorithm based on NTT increases with $d$, while the gap between this and our protocol’s local decryption time also grows, corresponding to the factor $1/\log{d}$ in the efficiency ratio $\alpha$. Overall, we achieve a speed-up of $43\%\sim 67\%$ under $\lambda=128$, and $14\%\sim 67\%$ under $\lambda=256$.
@@ VI-B Space Reduction
We observe that the OtsDec protocol helps save the memory needed for decryption. For the original decryption, its minimum space consumption commonly consists of three components: secret key $s\in R_{q}$, ciphertext $(u,v)\in R_{q}^{2}$, and parameters for performing $INTT$. $(u,v)$ takes up $2\ell d$ bits, where $\ell$ is the number of bits for storing one coefficient modulo $q$. Although $s$ is usually sampled from some distribution with very small deviation, it is the point-value form $\hat{s}\leftarrow INTT(s)$ instead of $s$ itself that should be stored. $\hat{s}$ requires $\ell d$ bits since the components of $\hat{s}$ is not necessarily small. The parameters for $INTT$ are actually powers of some root of unity in $R_{q}$ and also need $\ell d$ bits of memory. Such space composition can be relatively large due to large $d$ and $q$. Table 6 demonstrates the space consumption for original decryption implemented by the Lattigo library with $d=2^{15}$ and any $\log{q}\leq 64$.
The local decryption of OtsDec requires no invocation of $\hat{s}$ and NTT-related parameters, but only the sparse unblinding factor $t$ and the partially decrypted ciphertext $(\tilde{u},v)$. This saves up to nearly $50\%$ of space consumption since $t$ requires only $2\ell h$ bits for storing its non-zero coefficients and the corresponding indices, where $h$ may be much smaller than $d$. Table 5 demonstrates the space consumption for OtsDec’s local decryption in the direct connection mode with $d=2^{15}$ and any $\log{q}\leq 64$.
Figure 6: Original decryption space consumption with $d=2^{15}$ and $\log{q}\leq 64$
Figure 7: Outsourcing decryption space consumption with $d=2^{15}$ and $\log{q}\leq 64$
@@ VII Conclusion & Further Discussion
In this paper, we have presented an outsourced decryption protocol for RLWE-based homomorphic encryption schemes to alleviate clients’ local decryption burden in a cloud computing setting. We have demonstrated that the protocol preserves certain security under typical attacks and exhibits a significant speed-up and space reduction on client-side decryption. Nevertheless, there are several topics for further discussion. The security of this protocol could be further investigated. For example, subring and subfield attacks [53, 54, 55, 62] compromise NTRU in polynomial time when the coefficient modulus is exponentially high relative to the polynomial degree, but with the secret vector occupying a much smaller space. Thus it is uncertain how they may affect OtsDec’s security. Moreover, the proposed protocol is secure only if the RLWE and NTRU assumptions hold when both instances are published under the same secret. Therefore exploring the actual effect of such behavior is also required in future work.
@@ Acknowledgment
This work is supported by the Key Research and Development Program of Shandong Province, China (Grant No. 2022CXGC020102), the Science and Technology Small and Medium Enterprises (SMEs) Innovation Capacity Improvement Project of Shandong Province & Jinan City, China (Grant No. 2022TSGC2048), Haiyou Famous Experts - Industry Leading Talent Innovation Team Project, the Major Key Project of PCL (Grant No. PCL2022A03), and the National Natural Science Foundation of China under Grant(NO.62072247).

### v2

@@ V Security Analysis
We initially present formal proof of the security of the proposed outsourced decryption in Theorem V.1, reducing the protocol’s security to the RLWE and NTRU assumptions. This stems from the fact that our blinded secret key can be viewed as a $(\gamma_{s},\gamma_{t},q)$-NTRU instance given in Definition V.1. Recovering the secret key from the blinded one corresponds to solving the $(\gamma_{s},\gamma_{t},q)$-search NTRU problem in Definition V.2.
@@
The outsourced decryption is semantically secure, provided that the RLWE and NTRU-search assumptions hold where their instances are public together under the same secret.
@@
    • $pk,s\leftarrow\text{HE.KeyGen}(1^{\lambda})$,
    • $(t,t^{-1})\leftarrow\Pi_{\text{SkBdKeyGen}}(1^{\lambda})$,
    • $\tilde{s}\leftarrow\Pi_{\text{SkBd}}(s,t^{-1})$.
@@
    $\mathcal{U}_{H}^{\prime}$ computes $c\leftarrow\text{HE.Enc}(pk,m)$ where $m$ is sampled uniformly from the plaintext space.
@@
[56] A $(\gamma_{s},\gamma_{t},q)$-NTRU instance is $r\in R_{q}=Z_{q}[X]/(X^{n}+1)$ such that (i) $r=s\cdot t^{-1}\mod{q}$ and (ii) $||s||_{2}\leq\frac{\sqrt{q}}{\gamma_{s}},||t||_{2}\leq\frac{\sqrt{q}}{\gamma_{t}}$
@@
[56] The $(\gamma_{s},\gamma_{t},q)$-search NTRU problem asks, given a $(\gamma_{s},\gamma_{t},q)$-NTRU instance $r$, to recover $(s,t)\in R^{2}$ such that (i) $r=s\cdot t^{-1}\mod{q}$ and (ii) $||s||_{2}\leq\frac{\sqrt{q}}{\gamma_{s}},||t||_{2}\leq\frac{\sqrt{q}}{\gamma_{t}}$
@@ V-A Attacks on the secret key blinding technique
Given a blinded private key $\tilde{s}=st^{-1}$, along with public parameters $\{d,q,h\}$ for the blinding algorithm, the adversary’s goal is to recover either $s$ or $t$. $h$ is not necessarily a public parameter, but we assume it for the simplicity of security discussion. We also denote the set of all elements that $s$ can take as $S_{sk}$ and the security parameter as $\lambda$. Moreover, $\chi_{sk}$ is fixed as the ternary distribution over $R=\mathbb{Z}[X]/X^{d}+1$, which is a common choice for RLWE-based HE scheme [20].
@@
TABLE I: Lattice Reduction for $L(\tilde{s},\alpha)$ with $h=5$ and $\log{q}=33$
@@
The probability for us to guess a correct set of $J$ for $t$ or any of its shifted versions is computed with the following equation proposed by Silverman [59]:
@@
Taking the result of Equation (10) as $p$, the expected time for a zero-forced attack with specific $r$ to find $\tau$ or any of its shifting versions in some $L(\alpha,\tilde{s})^{ZF}$ is computed by $1/p\cdot\text{oprs}^{ZF}$, where $\text{oprs}^{ZF}$ denotes the number of operations required for reducing one $L(\alpha,\tilde{s})^{ZF}$ instance. Table 3 demonstrates the probability $p$ and the zero-forced attack complexity as $r$ increases with different choices of $h$. The complexity of the attack corresponds to the number of elementary operations in one BKZ iteration[12, 65]: $\log{(\text{oprs})}=poly(\beta)+\log{dim}+7$, where $poly(\beta)=0.00405892\beta^{2}-0.337913\beta+34.9018$.
@@ V-B Security Parameterization
Based on the attacks discussed above, we can parameterize the security of the outsourced decryption. We consider three security parameter cases: $\lambda=128,192,256$. For each $\lambda$, we discuss the required values of $h$ and $q$ when $d=13,14,15$, and $16$. During the parameterization, $h$ is selected such that a zero-forced attack requires a complexity no smaller than $2^{\lambda}$ elementary operations. For a given $h$, we fix the sampling method of $t$ as $t\leftarrow\Pi_{\text{SkBdKeyGen}}(R_{q},h_{1})\cdot(\Pi_{\text{SkBdKeyGen}}(R_{q_{2}},h_{2}))\mod{q}$, where $h_{1}\cdot h_{2}-\text{min}(h_{1},h_{2})\geq h$. The sampling space determined by $q$ and $h$ should be large enough to guard against brute-force attacks. For adversaries enumerating the $t$ using the sampling method above, $q$ and $h$ should satisfy the following equation:
@@
TABLE III: parameterization for the outsourced decryption protocol
@@ VI Implementation
After discussing the security of the outsourced decryption protocol, we move on to the implementation and performance evaluation.
We implement all the proposed algorithms with the Lattigo library V4.1.0 [14], where we use the CKKS scheme it supports as the underlying RLWE-based HE scheme. Our experiments are conducted on a machine equipped with an i7-13700K (3.40 GHz) processor and 64GB of memory.
To evaluate the protocol’s performance on practical RLWE-based HE parameter sets, we compare the local decryption time of the original decryption and our outsourced method under parameter sets with $d={13,14,15,16}$ (see Figure 4 and 5 for comparison results). One may refer to the FHE parameterization guidance proposed by Bossuat et al.[20] for the corresponding security level under the RLWE assumption. Note that for any fixed $d$, the security of the RLWE assumption increases as the modulus $q$ decreases. The minimum required $q$ in Table III is sufficient to exceed 256-bit security under the RLWE assumption. Thus, in terms of parameter selection, the outsourced decryption and the underlying HE scheme exhibit no explicit conflict.
Figure 4: Local decryption time with $\log{q}\leq 64$
Figure 5: Local decryption time with fixed $d=2^{15}$
@@ VI-A Efficiency
It can be observed that the differences between the curves for various $h$ are minimal, as in the sampling method of $t$, with $h_{1}$ fixed at $6$, $h_{2}$ only needs to be within the range $[3,8]$ to meet all three security categories. On the other hand, the time taken by the original decryption algorithm based on NTT increases with $d$, while the gap between this and our protocol’s local decryption time also grows, corresponding to the factor $1/\log{d}$ in the efficiency ratio. Overall, we achieve a speed-up of $43\%\sim 67\%$ under $\lambda=128$, and $14\%\sim 67\%$ under $\lambda=256$.
@@ VI-B Space Reduction
We observe that the proposed protocol helps save the memory needed for decryption. For the original decryption, its minimum space consumption commonly consists of three components: secret key $s\in R_{q}$, ciphertext $(u,v)\in R_{q}^{2}$, and parameters for performing $INTT$. $(u,v)$ takes up $2\ell d$ bits, where $\ell$ is the number of bits for storing one coefficient modulo $q$. Although $s$ is usually sampled from some distribution with very small deviation, it is the point-value form $\hat{s}\leftarrow INTT(s)$ instead of $s$ itself that should be stored. $\hat{s}$ requires $\ell d$ bits since the components of $\hat{s}$ is not necessarily small. The parameters for $INTT$ are actually powers of some root of unity in $R_{q}$ and also need $\ell d$ bits of memory. Such space composition can be relatively large due to large $d$ and $q$. Table 6 demonstrates the space consumption for original decryption implemented by the Lattigo library with $d=2^{15}$.
On the other hand, $\Pi_{\text{LocDec}}$ requires no invocation of $\hat{s}$ and NTT-related parameters, but only the sparse unblinding factor $t$ and the blindly decrypted ciphertext $(\tilde{u},v)$. This saves up to nearly $50\%$ of space consumption since $t$ requires only $2\ell(h_{1}+h_{2})$ bits for storing its non-zero coefficients and the corresponding indices, where $h<h_{1}*h_{2}-\text{min}(h_{1},h_{2})$ and $h$ is much smaller than $d$. Table 5 demonstrates the space consumption for the local decryption with $d=2^{15}$ and various $q$.
Figure 6: Original decryption space consumption with $d=2^{15}$
Figure 7: Outsourced decryption space consumption with $d=2^{15}$
@@ VII Conclusion & Further Discussion
In this paper, we have presented an outsourced decryption protocol for RLWE-based homomorphic encryption schemes to alleviate clients’ local decryption burden in a cloud computing setting. We have demonstrated that the protocol preserves certain security under typical attacks and exhibits a significant speed-up and space reduction on client-side decryption. Nevertheless, there are several topics for further discussion.
The security of this protocol could be further investigated. For example, subring and subfield attacks [60, 61, 62, 69] compromise NTRU in polynomial time when the coefficient modulus is exponentially high relative to the polynomial degree, but with the secret vector occupying a much smaller space. Thus it is uncertain how they may affect the security of the proposed secret key blinding technique. Moreover, the proposed protocol is secure only if the RLWE and NTRU assumptions hold when both instances are published under the same secret. Therefore exploring the actual effect of such behavior is also required in future work.
Moreover, a verification method can be provided for the protocol to adapt it to more variable scenarios. One possible probabilistic verification method is to use polynomial evaluation. Client can prepare some random points, and verify correctness by comparing the evaluation results returned by the cloud with those of the original decryption algorithm at these points. Further discussion is needed to determine the feasibility of this approach.
@@ Acknowledgment
This work is supported by the Key Research and Development Program of Shandong Province, China (Grant No. 2022CXGC020102), the Science and Technology Small and Medium Enterprises (SMEs) Innovation Capacity Improvement Project of Shandong Province & Jinan City, China (Grant No. 2022TSGC2048), Haiyou Famous Experts - Industry Leading Talent Innovation Team Project, and the National Natural Science Foundation of China under Grant(NO.62072247).